Given algorithm- procedure factorial (n: nonnegative integer) if n = 0 then return 1 else return n*factorial(n-1) {output is n!} Trace the above algorithm when it is given n = 7 as input. That is, show all steps used by above algorithm to find 7!
Answer #1

Trace:

factorial(1) = 1
factorial(2) = 2 * factorial(1) = 2 * 1 = 2
factorial(3) = 3 * factorial(2) = 3 * 2 = 6
factorial(4) = 4 * factorial(3) = 4 * 6 = 24
factorial(5) = 5 * factorial(4) = 5 * 24 = 120
factorial(6) = 6 * factorial(5) = 6 * 120 = 720
factorial(7) = 7 * factorial(6) = 7 * 720 = 5040